# Multimodal atmospheric super-resolution with deep generative models

### v1

@@ Abstract
Score-based diffusion modeling is a generative machine learning algorithm that can be used to sample from complex distributions. They achieve this by learning a score function, i.e., the gradient of the log-probability density of the data, and reversing a noising process using the same. Once trained, score-based diffusion models not only generate new samples but also enable zero-shot conditioning of the generated samples on observed data. This promises a novel paradigm for data and model fusion, wherein the implicitly learned distributions of pretrained score-based diffusion models can be updated given the availability of online data in a Bayesian formulation. In this article, we apply such a concept to the super-resolution of a high-dimensional dynamical system, given the real-time availability of low-resolution and experimentally observed sparse sensor measurements from multimodal data. Additional analysis on how score-based sampling can be used for uncertainty estimates is also provided. Our experiments are performed for a super-resolution task that generates the ERA5 atmospheric dataset given sparse observations from a coarse-grained representation of the same and/or from unstructured experimental observations of the IGRA radiosonde dataset. We demonstrate accurate recovery of the high dimensional state given multiple sources of low-fidelity measurements. We also discover that the generative model can balance the influence of multiple dataset modalities during spatiotemporal reconstructions.
@@ 1 Introduction
Given a dataset sampled from an unknown distribution, the goal of a generative model is to approximate this distribution with a model that allows sampling, inference, or density estimation [12]. Broadly, generative models fall into two categories: explicit density models, which define a likelihood function and aim to estimate it directly, and implicit density models, which focus on generating samples without requiring an explicit form of the likelihood. A state-of-the-art generative machine learning algorithm is the so-called score-based diffusion model [16, 17, 18, 19] – a focus of this study. In this approach, one starts with data (such as images) and gradually adds noise over time through a forward stochastic process—typically a diffusion process—until the data becomes nearly indistinguishable from pure Gaussian noise. This forward process is mathematically described using stochastic differential equations (SDEs), and it defines a continuous trajectory from clean data to random noise [20]. It is important to note that score-based models learn to approximate, using a neural network, the score function—the gradient of the log-probability density of the data at various noise levels. This score function tells the model how to adjust a noisy sample to make it more likely under the data distribution. Finally, to generate new data, the model starts with a sample of pure noise and simulates the reverse-time SDE using the learned score function, effectively “denoising" the sample step-by-step. In essence, a trained score-based diffusion model implicitly learns an unknown distribution given collected samples. Furthermore, score-based models can also be used to adapt samples given sparse observations of the generated samples. This is achieved using ‘zero-shot’ score matching [21, 22], where a pretrained score-based diffusion model is conditioned on new, sparse observations—without retraining the model. The core idea is to modify the sampling process using observations (constraints) to guide the generation, effectively sampling from a conditional distribution.
In this article, we develop score-based diffusion models for generating realistic states of a high-dimensional multiscale system (the atmosphere) from sampled noise. Subsequently, we utilize our pretrained model for rapidly generating samples that assimilate observations from coarse grids or from different sensor measurements. We outline specific contributions in the following
@@
    We train a score-based diffusion model using ERA5 reanalysis data [23] to generate realistic samples corresponding to the global atmosphere for 69 variables on a 1.4-degree resolution grid.
@@
  3. 3.
    We perform zero-shot score-matching using partial observations of different datasets to demonstrate a multifidelity super-resolution task with quantified uncertainty.
Our results demonstrate that score-based generative models can be used to generate high-dimensional states of dynamical systems (i.e., in a super-resolution task) from a variety of data sources without retraining using zero-shot score-matching. Moreover, we also demonstrate efficient ‘data-fusion’ where zero-shot samples can balance the influence of multiple sources of data that are partially observed. This allows for the utilization of different sources of data available at different spatial resolutions for posterior updates during the generative sampling process.
@@ 1.1 Related work
In recent studies, deep learning-based models have provided exciting results in the state-reconstruction of high-dimensional multiscale dynamical systems [24, 25, 26]. For example, in applications related to the reconstruction of turbulent flows and for magnetic resonance imaging data, customized physics-informed deep learning methods have led to the remarkable recovery of fine-scaled features from coarse-grained observations [27, 28]. Most work in this emergent area of data-driven super-resolution has relied on structured grid representations of both the coarse and fine representations of the state which has allowed for the use of convolutional neural network techniques [29, 30]. For unstructured representations, graph neural network applications have also been used with significant success [31]. Methods that bridge convolutional architectures and arbitrary locations of sparse observations using Voronoi tessellation have also led to impressive state reconstructions [32]. However, a significant majority of these super-resolution demonstrations possess two limitations: first, they frequently rely on coarse-grid representations extracted from the fine state (i.e., they are obtained from the same dataset). Second, the proposed models need retraining when the source or representation of coarse-grid data is significantly changed for example when it is obtained from a different dataset. Finally, since super-resolution is an inverse problem, several deterministic techniques face limitations in deployments without some form of uncertainty quantification.
@@
In this work, we utilized generative deep learning using EDM models to enable the super-resolution of high-dimensional states from dynamical systems from arbitrary coarse-grid observations without retraining. We also jointly provide UQ without the aforementioned restrictive assumptions through sampling from the generative process. Before proceeding, we review related work that has leveraged generative models for super-resolution of high-dimensional dynamical systems.
@@ 2.1 Forward and Reverse Diffusion Processes
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma))=-\frac{\mathbb{E}_{\mathbf{x}_{0}|\mathbf{x}(\sigma)}[\mathbf{x}(\sigma)-\mathbf{x}_{0}]}{\sigma^{2}}$ (3)
A neural network, $D_{\theta}$, is trained to approximate this score. If the network is trained to predict the clean data, $D_{\theta}(\mathbf{x}(\sigma);\sigma)\approx\mathbf{x}_{0}$, its output can be used to estimate the score:
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma);\sigma)\approx\frac{D_{\theta}(\mathbf{x}(\sigma);\sigma)-\mathbf{x}(\sigma)}{\sigma^{2}}$ (4)
Alternatively, if the network predicts the noise, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}(\sigma);\sigma)\approx\mathbf{n}$:
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma);\sigma)\approx-\frac{\boldsymbol{\epsilon}_{\theta}(\mathbf{x}(\sigma);\sigma)}{\sigma}$ (5)
@@ 2.2 EDM Formulation and Training
$F_{\theta}(\mathbf{x}(\sigma);\sigma)=c_{\text{skip}}(\sigma)\mathbf{x}(\sigma)+c_{\text{out}}(\sigma)D_{\theta}\left(\frac{\mathbf{x}(\sigma)}{c_{\text{in}}(\sigma)};c_{\text{noise}}(\sigma)\right)$ (6)
where $D_{\theta}$ is the neural network parameterized by $\theta$. The scaling factors ($c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$) are crucial for ensuring the network operates on inputs with approximately unit variance. We use the specific coefficient definitions from Table 1 in [41]. The training objective is a weighted mean squared error:
$L(\theta)=\mathbb{E}_{\mathbf{x}_{0},\mathbf{n},\sigma}\left[\lambda(\sigma)||\mathbf{x}_{0}-F_{\theta}(\mathbf{x}_{0}+\sigma\mathbf{n};\sigma)||^{2}\right]$ (7)
where $\sigma$ is sampled from a distribution $p(\sigma)$ and $\lambda(\sigma)$ is a weighting function, typically $\lambda(\sigma)=1/c_{\text{out}}(\sigma)^{2}$, that balances the loss contribution across different noise levels.
@@ 2.3 Sampling
Once trained, the model $F_{\theta}$ generates samples by numerically solving the reverse ODE. Starting from a noise sample $\mathbf{x}_{\sigma_{\text{max}}}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I})$, an ODE solver iteratively refines the sample across a sequence of decreasing noise levels $\sigma_{i}$. A first-order Euler step from $\mathbf{x}(\sigma_{i})$ to $\mathbf{x}(\sigma_{i+1})$ is:
@@
    Predict denoised data: $\hat{\mathbf{x}}_{0}=F_{\theta}(\mathbf{x}(\sigma_{i});\sigma_{i})$.
@@
    Calculate drift: $\mathbf{d}=(\mathbf{x}(\sigma_{i})-\hat{\mathbf{x}}_{0})/\sigma_{i}$.
@@
    Take step: $\mathbf{x}(\sigma_{i+1})=\mathbf{x}(\sigma_{i})+\mathbf{d}(\sigma_{i+1}-\sigma_{i})$.
More sophisticated samplers are often used to improve quality and efficiency [41, 73]. Several applications require guiding generation with a condition $\mathbf{y}$ to sample from $p_{\theta}(\mathbf{x}|\mathbf{y})$. This is often done by feeding an embedding of $\mathbf{y}$ into the network $D_{\theta}(\mathbf{x}(\sigma),\sigma,\mathbf{y})$. However, it requires the conditions to be known a priori for training.
2.4 Diffusion Posterior Sampling and Data Assimilation
@@
$\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|\mathbf{y})\approx\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ (8)
This formulation is structurally analogous to guidance methods like CFG [74], where the prior score (from the unconditional model) is modified by a guidance term (the difference between conditional and unconditional scores). Here, the guidance term is the gradient of the log-likelihood. A key challenge is estimating this likelihood score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$.
For a differentiable measurement model $M$ and Gaussian observation noise $\Sigma_{y}$, DPS [46] approximates the likelihood $p(\mathbf{y}|\mathbf{x}_{t})$ by first estimating the clean data $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)$ from the noisy sample $\mathbf{x}_{t}$ and then evaluating the likelihood at this estimate:
$p(\mathbf{y}|\mathbf{x}_{t})\approx\mathcal{N}(\mathbf{y}|M(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)),\Sigma_{y})$ (9)
@@
$\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\approx\frac{\mathbf{x}_{t}+\sigma_{t}^{2}\mathbf{s}_{\phi}(\mathbf{x}_{t},t)}{\mu_{t}}$ (10)
where $\mathbf{s}_{\phi}$ is the learned score model, and $\mu_{t},\sigma_{t}$ are scaling factors. Score-based Data Assimilation (SDA) [76] refines this by accounting for the variance of the estimate $\hat{\mathbf{x}}_{0}$, yielding a more stable perturbed likelihood:
$p(\mathbf{y}|\mathbf{x}_{t})\approx\mathcal{N}\left(\mathbf{y}\mid M(\hat{\mathbf{x}}_{0}),\Sigma_{y}+\frac{\sigma_{t}^{2}}{(\mu_{t})^{2}}\mathbf{J}_{M}(\hat{\mathbf{x}}_{0})\mathbf{\Gamma}\mathbf{J}_{M}(\hat{\mathbf{x}}_{0})^{T}\right)$ (11)
where $\mathbf{J}_{M}$ is the Jacobian of the measurement function and $\mathbf{\Gamma}$ is the term depends on the eigen-decomposition of $\Sigma_{x}$. In practice, to simplify the approximation, the term $\mathbf{J}_{M}\mathbf{\Gamma}\mathbf{J}_{M}^{T}$ is replaced by a constant (diagonal) matrix (refer section 3.2 in [76]).
2.5 Diffusion Data Assimilation for EDM
In our work, we adapt the data assimilation process for the EDM framework. As the network $F_{\theta}$ is trained to predict $\mathbf{x}_{0}$, the score is given by Eq. 4. In the SDE convention used by EDM ($\mu_{t}=1$), substituting this into Tweedie’s formula (Eq. 10) gives a direct and intuitive estimate for the clean data:
$\hat{\mathbf{x}}_{0}=F_{\theta}(\mathbf{x}_{t};\sigma_{t})$ (12)
The likelihood score is then computed based on the SDA approximation (Eq. 11), simplified for the EDM case:
$\mathbf{s}_{l}(\mathbf{x}_{t},\sigma_{t};\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\propto\nabla_{\mathbf{x}_{t}}\frac{-||\mathbf{y}-M(F_{\theta}(\mathbf{x}_{t};\sigma_{t}))||^{2}}{\Sigma+\hat{\Gamma}\sigma_{t}^{2}}$ (13)
where $\Sigma$ is the variance of noise observation and $\hat{\Gamma}$ approximates $\mathbf{J}_{M}\mathbf{\Gamma}\mathbf{J}_{M}^{T}$. These hyperparameters are constants that are tuned based on the precision of measurements and desired accuracy. This likelihood gradient is then used to guide the standard EDM sampling process. The first-order update step becomes:

### v2

@@ Abstract
Diffusion models are a class of generative machine learning algorithms that can be used to sample from complex distributions. They achieve this by learning a score function, i.e., the gradient of the log-probability density of the data, and reversing a noising process using the same. Once trained, these diffusion models not only generate new samples but also enable zero-shot conditioning of the generated samples on observed data. This promises a novel paradigm for data and model fusion, wherein the implicitly learned distributions of pretrained diffusion models can be updated given the availability of online data in a Bayesian formulation. In this article, we apply such a concept to the super-resolution of a high-dimensional dynamical system, given the real-time availability of low-resolution and experimentally observed sparse measurements from multimodal data. Our experiments are performed for a super-resolution task that generates the ERA5 atmospheric dataset given sparse observations from a coarse-grained representation of the same and/or from unstructured experimental observations of the IGRA radiosonde dataset. We also perform a data fusion task that leverages predictions from a data-driven atmospheric emulator as well. We discover that the generative model can balance the influence of multiple dataset modalities during spatiotemporal state reconstructions. Additional analysis on how score-based sampling can be used for uncertainty estimates is also provided.
@@ 1 Introduction
Given a dataset sampled from an unknown distribution, the goal of a generative model is to approximate this distribution with a model that allows sampling, inference, or density estimation [12]. Broadly, generative models fall into two categories: explicit density models, which define a likelihood function and aim to estimate it directly, and implicit density models, which focus on generating samples without requiring an explicit form of the likelihood. A state-of-the-art generative machine learning algorithm is the so-called score-based diffusion model [16, 17, 18, 19] – a focus of this study. In this approach, one starts with data (such as images) and gradually adds noise over time through a forward stochastic process—typically a diffusion process—until the data becomes nearly indistinguishable from pure Gaussian noise. This forward process is mathematically described using stochastic differential equations (SDEs), and it defines a continuous trajectory from clean data to random noise [20]. It is important to note that score-based models learn to approximate, using a neural network, the score function—the gradient of the log-probability density of the data at various noise levels. This score function tells the model how to adjust a noisy sample to make it more likely under the data distribution. Finally, to generate new data, the model starts with a sample of pure noise and simulates the reverse-time SDE using the learned score function, effectively “denoising" the sample step-by-step. In essence, a trained diffusion model implicitly learns an unknown distribution given collected samples. Furthermore, these models can also be used to adapt samples given sparse observations of the generated samples. This is achieved using ‘zero-shot’ score matching [21, 22], where a pretrained score-based diffusion model is conditioned on new, sparse observations—without retraining the model. The core idea is to modify the sampling process using observations (constraints) to guide the generation, effectively sampling from a conditional distribution.
In this article, we develop a diffusion model for generating realistic states of a high-dimensional multiscale system (the atmosphere) from sampled noise. Subsequently, we utilize our pretrained model for rapidly generating samples that assimilate observations from coarse grids or from different sensor measurements. We outline specific contributions in the following
@@
    We train a diffusion model using ERA5 reanalysis data [23] to generate realistic samples corresponding to the global atmosphere for 13 variables on a 1.4-degree resolution grid.
@@
    We implement a stochastic sampler for our diffusion model that seamlessly performs multi-modal super-resolution and data fusion along with uncertainty quantification.
  3. 3.
@@
  4. 4.
    We perform zero-shot score-matching using partial observations of different datasets, including radiosonde observations and a data-driven atmospheric emulator, to demonstrate a multi-fidelity super-resolution task with quantified uncertainty.
Our results demonstrate that diffusion models can be used to generate high-dimensional states of dynamical systems (i.e., in a super-resolution task) from a variety of data sources without retraining using zero-shot score-matching. Moreover, we also demonstrate efficient ‘data-fusion’ where zero-shot samples can balance the influence of multiple sources of data that are partially observed. This allows for the utilization of different sources of data available at different spatial resolutions for posterior updates during the generative sampling process.
@@ 1.1 Related work
In recent studies, deep learning-based models have provided exciting results in the state-reconstruction of high-dimensional multiscale dynamical systems [24, 25, 26]. For example, in applications related to the reconstruction of turbulent flows and for magnetic resonance imaging data, customized physics-informed deep learning methods have led to the remarkable recovery of fine-scaled features from coarse-grained observations [27, 28]. Most work in this emergent area of data-driven super-resolution has relied on structured grid representations of both the coarse and fine representation of the state which has allowed for the use of convolutional neural network techniques [29, 30]. For unstructured representations, graph neural network applications have also been used with significant success [31]. Methods that bridge convolutional architectures and arbitrary locations of sparse observations using Voronoi tessellation have also led to impressive state reconstructions [32]. However, a significant majority of these super-resolution demonstrations possess two limitations: first, they frequently rely on coarse-grid representations extracted from the fine state (i.e., they are obtained from the same dataset). Second, the proposed models need retraining when the source or representation of coarse-grid data is significantly changed for example when it is obtained from a different dataset. Finally, since super-resolution is an inverse problem, several deterministic techniques face limitations in deployments without some form of uncertainty quantification.
@@
In this work, we utilize generative deep learning using EDM models to enable the super-resolution of high-dimensional states from dynamical systems from arbitrary coarse-grid observations without retraining. We also jointly provide UQ without the aforementioned restrictive assumptions through sampling from the generative process. Before proceeding, we review related work that has leveraged generative models for super-resolution of high-dimensional dynamical systems.
@@ 2.1 Forward and Reverse Diffusion Processes
$S(x(\sigma),\sigma)=\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma))=-\frac{\mathbb{E}_{\mathbf{x}_{0}|\mathbf{x}(\sigma)}[\mathbf{x}(\sigma)-\mathbf{x}_{0}]}{\sigma^{2}}$ (3)
A neural network, $D_{\theta}$, that estimates the clean data for different noise levels is used to indirectly approximate this score. If the network is trained to predict the clean data, $D_{\theta}(\mathbf{x}(\sigma);\sigma)\approx\mathbf{x}_{0}$, its output can be used to estimate the score:
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma);\sigma)\approx\frac{D_{\theta}(\mathbf{x}(\sigma);\sigma)-\mathbf{x}(\sigma)}{\sigma^{2}}=\mathbf{s}_{\theta}(x(\sigma),\sigma)$ (4)
@@ 2.2 EDM Formulation and Training
$D_{\theta}(\mathbf{x}(\sigma);\sigma)=c_{\text{skip}}(\sigma)\mathbf{x}(\sigma)+c_{\text{out}}(\sigma)F_{\theta}\left(\frac{\mathbf{x}(\sigma)}{c_{\text{in}}(\sigma)};c_{\text{noise}}(\sigma)\right)$ (5)
where $F_{\theta}$ is the neural network parameterized by $\theta$. The scaling factors ($c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$) are crucial for ensuring the network operates on inputs with approximately unit variance. We use the specific coefficient definitions from Table 1 in [41]. The training objective is a weighted mean squared error:
[…]
where $\sigma$ is sampled from a distribution $p(\sigma)$ and $\lambda(\sigma)$ is a weighting function, typically $\lambda(\sigma)=1/c_{\text{out}}(\sigma)^{2}$, that balances the loss contribution across different noise levels. Following Karras et al. (2022) [41], we use $\sigma$ and $t$ interchangeably as $\sigma(t)=t$.
@@ 2.3 Sampling
Once trained, the EDM model generates samples by numerically solving the probability flow ODE [41],
$\frac{d\mathbf{x}}{dt}=\frac{D_{\theta}(\mathbf{x}(t);t)-\mathbf{x}(t)}{t}$ (7)
Starting from a noise sample $\mathbf{x}_{t_{\text{max}}}\sim\mathcal{N}(\mathbf{0},t_{\text{max}}^{2}\mathbf{I})$, an ODE solver iteratively refines the sample across a sequence of decreasing noise levels $t_{i}$. A first-order Euler step from $\mathbf{x}(t_{i})$ to $\mathbf{x}(t_{i+1})$ is:
@@
    Predict denoised data: $\hat{\mathbf{x}}_{0}=D_{\theta}(\mathbf{x}(t_{i});t_{i})$.
@@
    Calculate drift: $\mathbf{d}=(\mathbf{x}(t_{i})-\hat{\mathbf{x}}_{0})/t_{i}$.
@@
    Take step: $\mathbf{x}(t_{i+1})=\mathbf{x}(t_{i})+\mathbf{d}(t_{i+1}-t_{i})$.
More sophisticated samplers are often used to improve quality and efficiency [41, 73]. Several applications require guiding generation with a condition $\mathbf{y}$ to sample from $p_{\theta}(\mathbf{x}|\mathbf{y})$. This is often done by feeding an embedding of $\mathbf{y}$ into the network $D_{\theta}(\mathbf{x}(t),t,\mathbf{y})$. However, it requires the conditions to be known a priori for training.
2.4 Diffusion Posterior Sampling and Data Fusion
@@
$\nabla_{\mathbf{x}(t)}\log p_{\theta}(\mathbf{x}(t)|\mathbf{y})=\nabla_{\mathbf{x}(t)}\log p_{\theta}(\mathbf{x}(t))+\nabla_{\mathbf{x}(t)}\log p(\mathbf{y}|\mathbf{x}(t))$ (8)
This formulation is structurally analogous to guidance methods like CFG [74], where the prior score (from the unconditional model) is modified by a guidance term (the difference between conditional and unconditional scores). Here, the guidance term is the gradient of the log-likelihood. A key challenge is estimating this likelihood score $\nabla_{\mathbf{x}(t)}\log p(\mathbf{y}|\mathbf{x}(t))$.
For a differentiable measurement model $M$ and Gaussian observation noise with covariance $\Sigma_{y}$, DPS [46] approximates the likelihood $p(\mathbf{y}|\mathbf{x}(t))$ by first estimating the clean data $\hat{\mathbf{x}}_{0}(\mathbf{x}(t),t)$ from the noisy sample $\mathbf{x}(t)$ and then evaluating the likelihood at this estimate:
$p(\mathbf{y}|\mathbf{x}(t))\approx\mathcal{N}(\mathbf{y}|M(\hat{\mathbf{x}}_{0}(\mathbf{x}(t))),\Sigma_{y})$ (9)
@@
$\hat{\mathbf{x}}_{0}(\mathbf{x}(t))=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}(t)]\approx\frac{\mathbf{x}(t)+\sigma(t)^{2}\mathbf{s}_{\theta}(\mathbf{x}(t),t)}{\mu(t)}$ (10)
where $\mathbf{s}_{\theta}$ is the learned score model, and $p(x(t)|x_{0})=\mathcal{N}(x(t)|\mu(t)x_{0},\sigma(t)^{2}I)$. Score-based Data Assimilation (SDA) [76] refines this by accounting for the variance of the estimate $\hat{\mathbf{x}}_{0}$, yielding a more stable perturbed likelihood:
$p(\mathbf{y}|\mathbf{x}(t))\approx\mathcal{N}\left(\mathbf{y}\mid M(\hat{\mathbf{x}}_{0}),\Sigma_{y}+\frac{\sigma(t)^{2}}{(\mu(t))^{2}}\mathbf{J}_{M}(\hat{\mathbf{x}}_{0})\mathbf{\Gamma}\mathbf{J}_{M}(\hat{\mathbf{x}}_{0})^{T}\right)$ (11)
where $\mathbf{J}_{M}$ is the Jacobian of the measurement function and $\mathbf{\Gamma}$ is the term that depends on the eigen-decomposition of the covariance of prior $p(x)$, given by $\Sigma_{x}$. In practice, to simplify the approximation, the term $\mathbf{J}_{M}\mathbf{\Gamma}\mathbf{J}_{M}^{T}$ is replaced by a constant (diagonal) matrix (refer section 3.2 in [76]).
2.5 Diffusion Data Fusion for EDM
In our work, we adapt the data fusion process for the EDM framework. As the network $F_{\theta}$ is trained to predict $\mathbf{x}_{0}$, the score is given by Eq. 4. In the SDE convention used by EDM ($\mu(t)=1$), substituting this into Tweedie’s formula (Eq. 10) gives a direct and intuitive estimate for the clean data:
$\hat{\mathbf{x}}_{0}=D_{\theta}(\mathbf{x}(t);\sigma(t))$ (12)
The likelihood score, $\mathbf{s}_{l}$ is then computed based on the SDA approximation (Eq. 11), simplified for the EDM case:
[…]
where $\Sigma_{y}$ is the variance of observation noise and $\hat{\Gamma}$ approximates $\mathbf{J}_{M}\mathbf{\Gamma}\mathbf{J}_{M}^{T}$. These hyperparameters can be approximated as constants that are tuned based on the precision of the measurements. In highly non-linear measurements, a trade-off is observed between the desired accuracy and the quality of the generated sample. This likelihood gradient is then used to guide the standard EDM sampling process. For sampling, a first-order update can be given at each discrete timestep (or noise level) $t_{i}$ by the following steps: